OpenStudy (anonymous):

u + 2 = 4 i already know that answer is 2, but you subtract 2 from each sides correct?

OpenStudy (anonymous):

Yup subtract 2 from each sides

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u +2 -2 = 4-2 u = 2

Parth (parthkohli):

Hint: this is the most obvious problem. ___ plus two = 4.

OpenStudy (amistre64):

you could subtract 4 from each side

OpenStudy (amistre64):

u + 2 = 4 -4 -4 ---------- u - 2 = 0